OpenStudy (anonymous):

1. If a rock is thrown vertically upward from the surface of Mars with velocity 15 m/s, its height after t seconds is h = 15t-1.86t^2 a. What is the velocity of the rock after 2s? b. What is the velocity of the rock when its height is 25m on its way up? On its way down?

hartnn (hartnn):

v = dh/dt so can you differentiate 'h' w.r.t 't' ?

hartnn (hartnn):

bdw, v = velocity

OpenStudy (anonymous):

oh okay, sorry for the late reply :)

hartnn (hartnn):

no problem :)

hartnn (hartnn):

so what u got as v ?

OpenStudy (anonymous):

i'm trying to solve it :)

hartnn (hartnn):

okk, you'll need this formula \((x^n)' = n x^{n-1}\)

OpenStudy (anonymous):

okay, sorry if it take me too long, i kinda forgot about solving this :(

hartnn (hartnn):

hey, no problem, take your time :) tell me if you find it difficult, i will try to explain it :)

OpenStudy (anonymous):

ok, so i got \[\frac{ d(h) }{ d(t) }= 15-3.72t\]

hartnn (hartnn):

yes! thats your equation for velocity! v = 1.5 -3.72t for part a) you need v at time t = 2 so, just plug in t = 2 in that equation! :)

OpenStudy (anonymous):

ohh :)

OpenStudy (anonymous):

so, its 7.56 m/s? :)

hartnn (hartnn):

v(2) = 1.5 - 3.72(2) = ... ? its not 7.56

OpenStudy (anonymous):

oh no, sorry, 11.16 :)

OpenStudy (anonymous):

i pressed the wrong button on my sci cal sorry

OpenStudy (anonymous):

oh no, wait ... i thought its 15? and not 1.5? :O

hartnn (hartnn):

no, my error :P its 15 not 1.5 7.56 was correct! sorry

OpenStudy (anonymous):

haha okay thankyou :)

hartnn (hartnn):

now for part b h = 25 is given, so from 15t-1.86t^2 = 25 can you find 2 values of 't' ? itsa quadratic in t

OpenStudy (anonymous):

ohh okay, i'll try :)

OpenStudy (anonymous):

uhm, can i ask? why was it equated to 25? i get it that it's the height and all but why was it 25? it thought its d(h)/d(t)? :)

hartnn (hartnn):

"its height is 25m" so h= 25 is given and the only equation you have a "h" is h = 15t-1.86t^2 so we need to use it to find the 't'

OpenStudy (anonymous):

ohh, i get it now :)

OpenStudy (anonymous):

okay, so i got x1=5.71 and x2=2.35 :)

hartnn (hartnn):

yes, those are correct, now you need the velocity at those times.

hartnn (hartnn):

when the rock is 25m height on its way up, it took 2.35 seconds so to get the velocity when its at height 25 on its way up, you will put t= 2.35 in the equation of velocity.

hartnn (hartnn):

when the rock is 25m height on its way down, it took 5.71 seconds so to get the velocity when its at height 25 on its way down, you will put t= 5.71 in the equation of velocity

OpenStudy (anonymous):

the v = d/t formula right?

hartnn (hartnn):

yes,the equation you got for velocity v = 15-3.72 t

OpenStudy (anonymous):

okay, so how do you know that 2.35 is for its way up and the other for its way down? :)

hartnn (hartnn):

that is logic, since the rock is "thrown vertically up" it will reach the height of 25 m on its way up first, (so 2.35 sec ) then it still goes further up and at one point it will stop, then it will start coming down due to gravity and will again reach the height of 25m on its way down. since coming down took more time, we take 5.71 as t for its way down

OpenStudy (anonymous):

oh haha, thanks :)

OpenStudy (anonymous):

ok so I got 6.258 m/s for 2.35s, and -6.2412 m/s for 5.71s :)

hartnn (hartnn):

both are absolutely correct! :)

OpenStudy (anonymous):

yay :) so it kinda shows now in the answers, if it goes up, it has positive velocity and when it goes down, it has negative velocity. haha

hartnn (hartnn):

yes! correct lol , thats what we got too :)

OpenStudy (anonymous):

thankyou :) can i as like 3 more questions? haha its all word problems and i find them difficult esp. when they're on derivatives :(